OpenStudy (anonymous):

x^(sqrt(log(x)))=10^8

OpenStudy (anonymous):

\[x^{\sqrt{\log(x)}}=10^{8}\]

OpenStudy (freckles):

I would start by taking log( ) of both sides

OpenStudy (anonymous):

\[x ^{\sqrt{\log(x)}}=10^8\] \[\log \left( x ^{\sqrt{logx}} \right)=\log \left( 10^8 \right)\] \[\sqrt{logx}logx=8\log10\] \[\left( logx \right)^{\frac{ 3 }{ 2 }}=8*1=8\] \[\log x=8^{\frac{ 2 }{ 3 }}=\left( 2^3 \right)^\frac{ 2 }{ 3 }=2^2=4,x=10^4\]

OpenStudy (loser66):

perfect!

OpenStudy (anonymous):

thanks guys